OpenStudy (anonymous):

can you still use the foil method even tho there is a plus sign? (-1+6i)+(-4+2i)

OpenStudy (mertsj):

NO!!

OpenStudy (mertsj):

Add the like terms.

OpenStudy (jamesj):

No. Absolutely not. (a+b)(c+d) = ac + bd + ad + bc But (a+b) + (c+d) = a + b + c + d

OpenStudy (mertsj):

Rearrange the 4 terms.

OpenStudy (mertsj):

=1+(-4) +6i+2i

OpenStudy (jamesj):

(-1+6i)+(-4+2i) = (-1 + -4) + (6i + 2i) = ...

OpenStudy (mertsj):

Add 1 and -4 then add 6i and 2i

OpenStudy (mertsj):

You will get -3 + 8i

OpenStudy (anonymous):

-5+8i thanks guys

OpenStudy (jamesj):

Mertsj, you've dropped the minus sign in front of the 1.

OpenStudy (mertsj):

Whoops!!I It's -1 so you will get -5+8i

OpenStudy (mertsj):

Thanks James. My old eyes don't see so well anymore.

OpenStudy (anonymous):

FOIL means multiply, not add, sorry ...